OpenStudy (anonymous):

Which correctly describes how the graph of the inequality 4x - 2y > 20 is shaded? Above the solid line Below the solid line Above the dashed line Below the dashed line

OpenStudy (turingtest):

first, solve the inequality for y

OpenStudy (anonymous):

I'm not to sure on how to exactly do that. @TuringTest

OpenStudy (turingtest):

I would start by adding 2y to each side

OpenStudy (anonymous):

Okay so I add the 2y to what side exactly

OpenStudy (turingtest):

to both sides\[\begin{array}4x& - 2y& >& 20\\&+2y&&&+2y\\\hline\end{array}\]

OpenStudy (anonymous):

How would I solve this ?

OpenStudy (turingtest):

well on the left you are adding 2y to -2y what is 2y-2y ?

OpenStudy (anonymous):

2y-2y =y

OpenStudy (turingtest):

no\[2y-y=y\]

OpenStudy (anonymous):

4y

OpenStudy (turingtest):

\[2y+2y=4y\]

OpenStudy (turingtest):

you just add or subtract the coefficients\[\color{red}2y+\color{red}2y=\color{red}4y\]

OpenStudy (turingtest):

for example\[\color{red}2y+\color{red}3y=\color{red}5y\\\color{red}5y-\color{red}2y=\color{red}3y\\\color{red}2y-\color{red}1y=\color{red}1y\]so what is \[\color{red}2y-\color{red}2y\]

OpenStudy (anonymous):

-4y

OpenStudy (anonymous):

you there?

OpenStudy (turingtest):

no, that would be\[-\color{red}2y-\color{red}2y=-\color{red}4y\]

OpenStudy (turingtest):

just look at the coefficients (the red numbers)\[\color{red}2y-\color{red}2y\\\color{red}2~~-\color{red}2\]

OpenStudy (anonymous):

0

OpenStudy (anonymous):

So, would it be, Below the Dashed Line ?

OpenStudy (anonymous):

Someone please Help!

Mehek (mehek14):

shade below the line if y is less than the other side

OpenStudy (anonymous):

Okay, so I already know that it is below the line,m but is it a dashed line?

Mehek (mehek14):

is it just < or >

OpenStudy (anonymous):

>

Mehek (mehek14):

so dashed line

Mehek (mehek14):

if it was ≥ then a solid line

OpenStudy (anonymous):

oh okay, well thanks! Can you help me with some more please?

Mehek (mehek14):

not really i gtg

OpenStudy (anonymous):

okay but thanks!

Mehek (mehek14):

yw :)